OpenStudy (anonymous):

find the slope and y-intercept 5x-10y=-8

OpenStudy (♪chibiterasu):

First we have to convert this into slope intercept form: y=mx + b m is the slope and b is the y intercept. First, we have to subtract 5x from both sides like this: 5x - 5x - 10y = -8 - 5x -10y = -5x - 8 Now, divide -10 on both sides: -10y/-10 = -5x - 8/-10 y = -5x + 0.8 -5 is the slope and 0.8 is the y-intercept

OpenStudy (anonymous):

thank you very much!!! :)

OpenStudy (♪chibiterasu):

Your welcome c: